# Meta-learning-based percussion transcription and $t\bar{a}la$ identification from low-resource audio

### v2

@@ V-C1 MAML for ADT
For evaluation, the model trained during meta-testing is directly tested on the remaining portions of each dataset to compare its performance against the baseline approach. The MAML framework follows the same experimental setup as in tabla stroke transcription. The state-of-the-art CRNN model from [15] is also trained and evaluated on the same meta-testing data to ensure a consistent and fair performance comparison.
V-C2 Transfer learning for ADT
In the first step, the CRNN model is trained on $D_{D4}$ and subsequently fine-tuned using the same classes and samples from the meta-test datasets $D_{D1}$, $D_{D2}$, and $D_{D3}$. The fine-tuned model is then evaluated on the remaining portions of each dataset, as detailed in Section V-C1. The training employs a batch size of 32 over 100 epochs, with a learning rate 0.001. This experimental setup thoroughly assesses the model’s transfer learning capabilities across different datasets.
VI Results and Discussions
@@
Approach $Bheem$ $Cha$ $Dheem$ $Dhin$ $Num$ $Ta$ $Tha$ $Tham$ $Thi$ $Thom$ Average
OTM 83.2 47.3 55.6 51.6 76.4 64.8 70 73.3 61.1 67.1 65.04
@@
VI-A Tabla stroke transcription (TST)
Table III shows the average cross-validation results of pre-training on $D_{M}$ using the three methods from Section IV-A1. Predictions are made by identifying and thresholding local peaks in the network’s output activations. The threshold is chosen to maximize the f-score on the validation set and then applied to the test set.
TABLE IV: Step-2: TST Performance on $D_{T1}$ (f1-score)
Approach $Other$ $Ta$ $Na$ $Din$ $Dhet$ $Dhin$ Average
@@ VI-A Tabla stroke transcription (TST)
PTM1+ML 87 71.2 76.1 89.9 92.1 71.6 81.32
TABLE V: Step-2: TST Performance on $D_{T2}$ (f1-score)
Approach B D RB RT Average
Drum-pretrained [6] 2.1 44 17.5 48.7 28.05
Best set of D, RT, RB [6]
81.5 83 63.6 86 78.52
@@
Retrained Rohit [5]
80.1 83.3 34.1 84.3 70.4
PTM2 + TL 57.9 59 35.6 65.6 54.52
PTM1 + TL 59.4 72.1 38.6 71 60.28
PTM2 + ML 81 79.6 63.7 82.9 76.8
PTM1 + ML 87.3 85.8 69.4 89.9 83.1
TABLE VI: Step-2: TST Performance on $D_{T3}$ (11 stroke classes) (f1-score)
Approach $Dha$ $Dhin$ $Tin$ $Na$ $Tun$ $Kat$ $Ta$ $Dhage$ $Tirkita$ $Dhi$ $Ti$ Average
@@
  • PTM1:Proposed Transcription Method-1, PTM2:Proposed Transcription Method-2, TL:Transfer Learning, ML: Meta Learning
TABLE VII: Step-2: TST Performance on $D_{T3}$ (with 5 stroke classes) (f1-score)
Approach $Other1$ $Dhin$ $Dhage$ $Tirkita$ $Dhi$ $Ti$ Average
PTM2+TL 21.6 38.3 33.8 26.9 15.4 18.2 25.70
PTM1+TL 35.8 43.4 45.7 48.1 34.6 22.3 38.31
PTM2+ML 50.4 61.9 57.1 59.8 45.6 37.9 52.11
PTM1+ML 61.4 68.8 75.7 79.7 47.6 43.9 62.85
TABLE VIII: Step-2: TST Performance on $D_{T3}$ (with 6 stroke classes) (f1-score)
Approach $Other2$ $Dha$ $Tin$ $Na$ $Tun$ $Kat$ $Ta$ Average
PTM2+TL 31.5 17.2 13 37.3 19.4 24.6 28.9 24.55
PTM1+TL 42.3 19.5 22.6 47.7 59.1 25.8 50.3 38.19
PTM2+ML 54.1 40.1 29.7 47.7 48.9 66.3 67.4 50.6
PTM1+ML 60.1 46.2 36.7 56.8 85.9 72.3 82.4 62.91
TABLE IX: $T\bar{a}la$ identification performance on $D_{T3}$
Method Accuracy (%) Time (ms)
@@
TABLE X: ADT performance on DTD - $D_{D1}$
(f1-score)
Approach KD SD HH Average
SOTA CRNN [15] 42.1 65.3 58.8 55.4
CRNN + TL 54.3 68.7 63.9 62.3
CRNN + ML 73.9 84.2 82.1 80.1
TABLE XI: ADT performance on DTP - $D_{D2}$ (f1-score)
Approach BD SD CL HH BE TT RD CY Average
SOTA
CRNN [15]
54.4 53.1 41.5 56.5 34.7 42.6 47.5 39.1 46.2
CRNN + TL 68.7 65.3 52.1 64.4 48.3 61.5 58.4 53.5 59.0
CRNN + ML 83.5 82.8 72.1 85.3 68.5 77.2 78 73.9 77.6
TABLE XII: ADT performance on DTD - $D_{D3}$ (f1-score)
Approach KD SD CHH HH TT RD CY OT TMB Average
SOTA CRNN [15] 44.8 46.3 38.7 33.1 27.4 39.3 32.6 24.5 14.4 33.4
CRNN + TL 54.3 70 63.5 60.2 57.5 62.7 54.1 44.8 33.4 55.6
CRNN + ML 73.1 91.6 84.4 82.3 70.9 81.6 74.5 67.3 59.5 76.1
TABLE XIII: ADT performance on DTM - $D_{D3}$ (f1-score)
Approach KD SD CHH HH TT RD CY OT TMB Average
SOTA CRNN [15] 41.7 38.2 29.8 30.3 24.5 36.2 29.4 18.6 8.6 28.6
CRNN + TL 52.7 66.1 61.7 57.8 53 60.3 50.1 40.9 31.8 52.7
CRNN + ML 71.4 87.8 80.1 79.8 68.6 78.5 64.9 61.3 54.2 71.8
TABLE XIV: ADT results Comparison
(Proposed vs. [36])
Approach % f1-score
Trained on SLAKH [36] 50.0
Trained on STAR [36] 62.0
Dynamic FSL [36] 70.0
CRNN + ML 71.8
The results demonstrate that the proposed methods outperform the existing OTM. The OTM, utilizing a CNN, only considers the spatial dimension of the Mel spectrogram. In contrast, the proposed CRNN method integrates both spatial and temporal dimensions, capturing the attack-decay-sustain-release (ADSR) temporal model of strokes more effectively. PTM1, which utilizes all stroke frames, outperforms PTM2, which relies on a single frame. PTM2’s reliance on a single frame may lead to missed stroke detections if the initial frame is undetected, whereas PTM1 remains robust by considering subsequent frames.
Table V shows meta-learning and transfer learning results using PTM1 and PTM2 on test data $D_{T1}$ from the same domain, demonstrating PTM1 with meta-learning outperforms PTM2 and others. Similarly, Table V presents results on test data $D_{T2}$ from a similar domain, with PTM1 again outperforming PTM2. Tables VI to VIII display results on vastly different domain test data $D_{T3}$, confirming PTM1 with meta-learning consistently outperforms PTM2 across different class variations.
Tables V-VIII demonstrate that the meta-learning consistently outperforms transfer learning. MAML learns an initialization for rapid task adaptation through meta-learning, while transfer learning directly uses pre-trained weights, which may not optimize performance for the target task. Specifically, Table V shows PTM1 with meta-learning outperforming four OTM settings using transfer learning and data augmentations, attributed to CNNs’ limited ability to capture temporal relationships. Tables VI-VIII indicate performance degradation on $D_{T3}$ due to frequent tabla stroke overlap with vocals and other instruments in concert data, complicating stroke onset detection. The number of stroke classes tested shows minimal impact on performance across these analyses.
@@ VI-C Automatic drum transcription (ADT)
The performance of the proposed Transfer Learning and MAML-based approaches for ADT tasks is summarized in Tables X-XIII. These tables present results on datasets representing all ADT task types--DTD, DTP, and DTM--specifically $D_{DD1}$, $D_{DD2}$, and $D_{DD3}$. The comparison is made against the state-of-the-art CRNN model [15]. Additionally, Table XIV highlights the comparative performance of the proposed MAML approach with existing methods on $D_{D3}$ (DTM).
The results demonstrate that the proposed MAML-based method consistently outperformed existing approaches utilizing minimal target data across all ADT task types. This indicates the effectiveness of the proposed approach in scenarios with limited data availability.
@@ VII Conclusion
This paper presents a novel approach leveraging Model-Agnostic Meta-Learning (MAML) for Tabla Stroke Transcription (TST) and Automatic Drum Transcription (ADT), addressing the challenge of limited annotated datasets. The proposed method effectively transcribes tabla strokes and drum sounds across diverse datasets, including scenarios with drum or tabla strokes alone, percussion instruments with melodic accompaniments, and the singer’s voice, ensuring comprehensive validation and demonstrating robustness and adaptability.
Following TST, two new $t\bar{a}la$ identification methods are introduced, contributing to rhythmic analysis in Hindustani music. Experimental results highlight the superiority of the MAML-based approach over existing methods, consistently achieving better performance with minimal labelled data.
Future work will focus on improving transcription accuracy for concert datasets with minimal labeled data, leveraging advanced music source separation techniques and active learning strategies.

### v3

@@ V-B1 MAML for ADT
Meta-training is performed using five stroke classes from $D_{D4}$, while meta-testing is conducted across three scenarios involving $D_{D1}$, $D_{D2}$, and $D_{D3}$. The specific stroke class distributions and audio durations for both meta-training and testing are summarized in Table II. For evaluation, the model trained during meta-testing is directly tested on the remaining portions of each dataset to compare its performance against the baseline approach. The MAML framework follows the same experimental setup as in the tabla stroke transcription.
V-B2 Baseline Models for ADT Comparison
For ADT baseline comparisons, we evaluated three approaches: a simple CRNN-based classifier [17], an NMF-based method following the setup described in [47], and a transfer learning baseline. NMF templates were derived from $D_{D1}$ and $D_{D2}$ and adapted across all datasets. To assess the impact of source separation, we applied Demucs [48] to separate mixed audio and extracted the drum track, on which all methods were re-evaluated. All experiments used the same set of stroke classes and train-test splits as in the MAML setup to ensure fair comparison.
@@ V-B2 Baseline Models for ADT Comparison
Approach $Bheem$ $Cha$ $Dheem$ $Dhin$ $Num$ $Ta$ $Tha$ $Tham$ $Thi$ $Thom$ Avg.
OTM 83.2 47.3 55.6 51.6 76.4 64.8 70.0 73.3 61.1 67.1 65.04
@@
TABLE IV: TST Performance on $D_{T1}$
          (f1-score)
Approach $Other$ $Ta$ $Na$ $Din$ $Dhet$ $Dhin$ Avg.
PTM2 28.2 30.7 26.3 31.4 29.5 30.3 29.40
PTM1 34.9 36.8 33.1 38.5 35.7 37.6 36.11
@@
PTM1+ML 87.0 71.2 76.1 89.9 92.1 71.6 81.32
  • TL: Transfer Learning, ML: Meta Learning
TABLE V: TST Performance on $D_{T2}$ (f1-score)
Approach B D RB RT Avg.
Drum-pretrained [7] 2.1 44.0 17.5 48.7 28.05
Best set of D, RT, RB [7] 81.5 83.0 63.6 86.0 78.52
@@
Retrained Rohit [6] 80.1 83.3 34.1 84.3 70.40
PTM2 36.1 37.2 23.5 40.7 34.38
PTM1 43.8 46.8 28.6 47.7 41.73
PTM2 + TL 57.9 59.0 35.6 65.6 54.52
PTM1 + TL 59.4 72.1 38.6 71.0 60.28
PTM2 + ML 81.0 79.6 63.7 82.9 76.80
PTM1 + ML 87.3 85.8 69.4 89.9 83.10
TABLE VI: TST Performance on $D_{T3}$ (f1-score)
Approach $Dha$ $Dhin$ $Tin$ $Na$ $Tun$ $Kat$ $Ta$ $Dhage$ $Tirkita$ $Dhi$ $Ti$ Avg.
PTM2 15.5 25.5 8.5 21.0 14.5 17.2 15.9 21.4 17.6 14.3 8.8 16.38
PTM1 22.7 31.0 13.5 27.4 23.5 20.2 22.8 24.2 24.6 19.3 10.2 21.76
@@
TABLE VII: $T\bar{a}la$ Identification Performance on $D_{T3}$
Method
Accuracy
Time
@@
VI Results and Discussions
VI-A Tabla stroke transcription (TST)
Table IV presents the average cross-validation results of pre-training on $D_{M}$ using the methods from Section III-A1. Predictions are obtained by thresholding local peaks in network activations, with thresholds tuned on the validation set. The proposed methods outperform OTM, which captures only frequency-domain features. In contrast, the CRNN leverages both spectral and temporal information, modelling the ADS characteristics of strokes. PTM1, which uses all frames within a stroke, outperforms PTM2, which relies solely on onset frames—likely affected by annotation inconsistencies. Leveraging temporal context, PTM1 offers improved robustness. Additionally, CRNN models perform well with single-channel input, indicating that OTM’s 3-channel configuration is not essential.
Tables IVVII show results for PTM1 and PTM2 under meta-learning across test sets with varying data distributions. PTM1 consistently outperforms PTM2 and other baselines, demonstrating the advantage of MAML’s rapid task adaptation over fixed representations. Table V highlights PTM1 outperforming four OTM settings, emphasizing CNN’s limitations in modeling temporal dependencies. Performance drop on $D_{T3}$ (Table VII) is due to stroke overlap with vocals and instruments, hindering onset detection.
The comparison with NMF-based and source separation methods is excluded for TST due to several limitations. Unlike ADT, TST lacks isolated instrument recordings, and using templates from the test dataset would create an unfair comparison. Additionally, no publicly available source separation model is designed for Indian classical music, and applying existing models like Demucs [48] leads to artifacts and domain mismatch, making them unsuitable for accurate stroke-level transcription.
@@ VI-C Automatic drum transcription (ADT)
Tables VIIIXI present the performance of the proposed MAML-based approach across all ADT task types—DTD, DTP, and DTM—compared with baselines including the state-of-the-art CRNN model [17], an NMF-based method [47], and transfer learning. For the DTM dataset, source separation [48] is applied, and results are reported accordingly. Table XII further compares MAML on $D_{D3}$ (DTM) against existing methods, using a 30ms collar for fair evaluation.
Among the baselines, the NMF-based method performs better than the supervised CRNN in low-data scenarios, showing greater robustness under limited supervision. However, its performance degrades on the DTM dataset due to interference from other instruments and vocals, which hampers reliable template construction in the absence of isolated tracks.
Overall, the MAML-based approach consistently outperforms all baselines across ADT tasks, demonstrating strong generalization and robustness in low-resource and real-world conditions.
TABLE VIII: ADT Performance on DTD - $D_{D1}$ (f1-score)
Approach KD SD HH Avg.
SOTA-CRNN [17] 42.1 65.3 58.8 55.4
PF-NMF [47] 68.4 72.5 73.0 71.3
CRNN + TL 54.3 68.7 63.9 62.3
CRNN + ML 73.9 84.2 82.1 80.1
TABLE IX: ADT Performance on DTP - $D_{D2}$ (f1-score)
Approach BD SD CL HH BE TT RD CY Avg.
SOTA-CRNN [17] 54.4 53.1 41.5 56.5 34.7 42.6 47.5 39.1 46.2
PF-NMF [47] 77.4 66.9 69.8 74.5 63.2 71.8 69.1 68.3 70.1
CRNN + TL 68.7 65.3 52.1 64.4 48.3 61.5 58.4 53.5 59.0
CRNN + ML 83.5 82.8 72.1 85.3 68.5 77.2 78.0 73.9 77.6
TABLE X: ADT Performance on DTD - $D_{D3}$ (f1-score)
Approach KD SD HH HH-O TT CY RD OT Avg.
SOTA-CRNN [17] 43.0 44.5 37.2 31.8 26.3 37.7 31.3 23.5 34.4
PF-NMF [47] 59.2 65.1 71.8 63.3 68.2 66.1 63.5 32.2 61.2
CRNN + TL 53.0 68.3 62.0 58.8 56.2 61.2 52.8 43.7 57.0
CRNN + ML 72.3 90.5 83.4 81.3 70.1 80.6 73.6 66.5 77.3
TABLE XI: ADT Performance on DTM - $D_{D3}$ (f1-score)
Approach KD SD HH HH-O TT CY RD OT Avg.
SOTA-CRNN [17] 40.4 37.0 28.9 29.3 23.7 35.1 28.5 18.0 30.1
PF-NMF [47] 41.3 39.7 35.6 33.8 30.4 36.9 32.6 24.9 34.4
CRNN + TL 51.3 64.3 60.0 56.2 51.5 58.6 48.7 39.8 53.8
CRNN + ML 70.0 86.1 78.5 78.2 67.3 76.9 63.6 60.1 72.6
SS + SOTA-CRNN 44.8 40.1 33.0 31.5 26.8 36.3 29.4 21.1 32.9
SS + PF-NMF 54.0 68.5 64.5 61.0 56.2 64.8 60.1 52.4 60.2
SS + CRNN + TL 53.1 67.7 62.3 57.5 53.4 60.2 51.6 42.9 56.1
SS + CRNN + ML 72.8 88.5 80.4 79.2 70.1 78.6 66.3 61.7 74.7
  • Note: Results use a 30ms collar; SS: Source Separation using Demucs [48]
TABLE XII: Comparison of ADT Performance on DTM - $D_{D3}$ with Existing Methods
Approach % f1-score
Proto-BL8-OFF [32] 67.0
Dense-BL8-OFF [32] 67.0
Proto-BL21-OFF [32] 69.0
Dense-BL21-OFF [32] 70.0
SOTA-CRNN [17] 30.1
SS + SOTA-CRNN 32.9
PF-NMF [47] 34.4
SS + PF-NMF 60.2
CRNN + TL 53.8
SS + CRNN + TL 56.1
CRNN + ML 72.6
SS + CRNN + ML 74.7
  • Note: Results use a 30ms collar and the same classes for fair comparison
                      with [32]; SS: Source Separation using Demucs [48]
@@ VII Conclusion
This paper proposes a novel Model-Agnostic Meta-Learning (MAML) based approach for Tabla Stroke Transcription (TST) and Automatic Drum Transcription (ADT), addressing challenges of limited annotated data and label heterogeneity. The method demonstrates effective transcription across diverse scenarios, including solo percussion, percussion with melodic accompaniments, and vocals, ensuring comprehensive validation and highlighting its robustness and adaptability.
Following TST, two novel $t\bar{a}la$ identification methods are introduced, advancing rhythmic analysis in Hindustani music. Experimental results show that the MAML-based approach consistently outperforms existing methods, even with limited labelled data.
Future work will focus on enhancing transcription accuracy for Indian music concert datasets with limited labelled data by leveraging advanced source separation techniques tailored to Indian music. We also aim to curate a larger and more diverse collection of polyphonic recordings to support training and evaluation in realistic conditions.